# OASIS: Open Agent Social Interaction Simulations with One Million Agents

### v4

@@ 1 Introduction
As the name suggests, ABMs consist of computational agents programmed to interact among themselves or with the environment in a realistic manner that is relevant to the complex system under study (Gilbert, 2019). Simulating agent behaviors is the key to designing ABMs. Traditionally, agent behaviors are programmed along measurable value (i.e., thresholds), which overlooks more complex aspects such as context-dependent behavioral changes. Recently, large language models (LLMs) have demonstrated remarkable capability to mimic human behaviors (Park et al., 2022; 2023; Zhou et al., 2023b; Wang et al., 2023; Gao et al., 2023; Mou et al., 2024; Chopra et al., 2024). LLM agents can engage in role-playing, i.e., impersonating human characters and taking part in a human-like interaction with other agents (Park et al., 2023; Zhou et al., 2023b), as well as taking a wide variety of actions ranging from simple decisions to more complex ones involving the tool use (Achiam et al., 2023). To develop and evaluate these LLM agents, researchers will need to move beyond standard benchmarks by defining social situations and distinct personas, as well as integrating these agents into simulated platforms or sandbox environments for more comprehensive testing and analysis (Park et al., 2023).
In the context of social media studies, popular social media platforms (i.e., X, Reddit) have drastically changed how people interact, exchange information, and form communities, making them crucial environments for studying modern social dynamics. They vary in how they design user interactions, henceforth termed action space, how they interact with users through algorithms and recommendation systems (RecSys), as well as how they connect with each other (Dynamic Network) For example, X facilitates a rapid exchange of views in real-time, and Reddit supports topic-based communities and emphasizes comment interaction. Consequently, users behave very differently across platforms, and as a result, several LLM-based ABM studies (see Table 1) have been proposed recently to study some aspects of social interactions on one of these platforms. Given the specific scenarios studied under these ABMs, pivoting them to study another domain remains tedious, which limits their usability to a larger social sciences community. Furthermore, these real-world social media contain millions of users. Simulating a large-scale ABM would allow for studies across multiple platforms, either individually or collectively, but it also introduces a wide range of engineering challenges. To this end, we propose OASIS, a collection of generalizable and scalable ABMs to simulate a wide variety of phenomena in various social media platforms.
@@
How OASIS works and why OASIS is generalizable? OASIS is built upon five foundational components, as shown in Figure 2, including the Environment Server, RecSys, Agent Module, Time Engine, and Scalable Inferencer. The Environment Server is initialized using generated or real-world data. It sends agents’ information, such as user descriptions and their relationships, along with posts, to the RecSys. The RecSys selects and pushes posts to agents through recommendation algorithms, determining the visibility of content for each agent. The Time Engine activates agents based on their temporal characteristics, enabling them to perform various actions such as commenting, posting, and interacting with other agents and the environment. These actions then update the environment’s state in real-time. All these components can be adapted easily to experiment with different social media platforms. For instance, by adjusting specific modules, switching from one platform, such as X, to another like Reddit is possible.
@@ 3.4.2 Herd Effect in Reddit
3.5 Misinformation Spreading in One Million Agents
Understanding the spread of true and false information is critical for combating misinformation and mitigating its influence on public opinion and behavior. To explore this, we designed a large-scale simulation of a social media platform resembling Twitter, with one million users. Within this environment, celebrity accounts with large followings were used to disseminate four distinct information sets, consisting of both factual and false narratives related to technology, health, education, and entertainment. Details are presented in G.0.1. We analyze the dynamics of how both true and false information spread among users in this large-scale simulation. In this experiment, the activation probability—which defines the likelihood of users sharing information—is set at 0.1 for core users and 0.01 for regular users. We visualize the changes in the number of posts related to both rumors and truthful information over time, as shown in Figure 9. Our findings indicate that rumors have a significantly broader reach than truthful information, which suggests that agents in the simulation tend to engage more with rumor posts that are novel and emotionally provocative (Vosoughi et al., 2018). We also use Neo4j to visualize the evolving social connections formed by agents during interactions in Figure 10, observing a tendency for users to form isolated groups over time. In this figure, the green dots represent agents, and the red lines indicate newly established relationships during the interaction process.
Figure 9: The relationship between the number of posts discussing various pairs of rumors and truths within the group over time.
Figure 10: The newly generated follow relationship map during the interaction process.
@@ 4 Ablation Study
4.1 Analysis of Efficiency for One Million Users
In this study, we report the runtime and GPU utilization for simulations at scales of one million, one hundred thousand, and ten thousand under a group polarization setting, as well as the number of tweets and comments added at each time step. For all scenarios, we use one A100 for RecSys and use multiple GPUs for LLM inference. We use vLLM (Kwon et al., 2023) to efficiently conduct LLM inference. As shown in Table 2, our algorithm can efficiently simulate large-scale user interactions. For instance, using five A100 GPUs, we can simulate the interactions of 100,000 users over 10 time steps within two days. Other scenarios’ efficiency analysis are presented in Appendix C.1.
Scale 1M 100K 10K
Hours per time step 18.0 3.0 0.2
GPUs (A100) 27.0 5.0 2.0
New Tweets per time step (K) 48.5 5.2 0.6
New Comments per time step (K) 97.1 9.0 0.9
Table 2: The analysis of efficiency for one million agents.

### v5

@@ 1 Introduction
As the name suggests, ABMs consist of computational agents programmed to interact among themselves or with the environment in a realistic manner that is relevant to the complex system under study (Gilbert, 2019). Simulating agent behaviors is the key to designing ABMs. Traditionally, agent behaviors are programmed along measurable value (i.e., thresholds), which overlooks more complex aspects such as context-dependent behavioral changes. Recently, large language models (LLMs) have demonstrated remarkable capability to mimic human behaviors (Park et al., 2022; 2023; 2024; Zhou et al., 2023b). LLM agents can engage in role-playing, i.e., impersonating human characters and taking part in a human-like interaction with other agents (Park et al., 2023; Zhou et al., 2023b), as well as taking a wide variety of actions ranging from simple decisions to more complex ones involving the tool use (Qin et al., 2023). To develop and evaluate these LLM agents, researchers will need to move beyond standard benchmarks by defining social situations and distinct personas, as well as integrating these agents into simulated platforms or sandbox environments for more comprehensive testing and analysis (Park et al., 2023).
In the context of social media studies, popular social media platforms (i.e., X, Reddit) have drastically changed how people interact, exchange information, and form communities, making them crucial environments for studying modern social dynamics Vosoughi et al. (2018). They vary in how they design user interactions, henceforth termed action space, how they interact with users through algorithms(Info Filter), as well as how they connect with each other (Dynamic Network) For example, X facilitates a rapid exchange of views in real-time, and Reddit supports topic-based communities and emphasizes comment interaction. Consequently, users behave very differently across platforms, and as a result, several LLM-based ABM studies (see Table 1) have been proposed recently to study some aspects of social interactions on one of these platforms. Given the specific scenarios studied under these ABMs, pivoting them to study another domain remains tedious, which limits their usability to a larger social sciences community. Furthermore, these real-world social media contain millions of users. Simulating a large-scale ABM would allow for studies across multiple platforms, either individually or collectively, but it also introduces a wide range of engineering challenges. To this end, we propose OASIS, a collection of generalizable and scalable ABMs to simulate a wide variety of phenomena in various social media platforms.
@@
How OASIS works and why OASIS is generalizable? OASIS is built upon five foundational components, as shown in Figure 2, including the Environment Server, Info Filter, Agent Module, Time Engine, and Scalable Inferencer. The Environment Server is initialized using generated or real-world data. It sends agents’ information, such as user descriptions and their relationships, along with posts, to the Info Filter. The Info filter selects and pushes posts to agents through filtering algorithms, determining the visibility of content for each agent. The Time Engine activates agents based on their temporal characteristics, enabling them to perform various actions such as commenting, posting, and interacting with other agents and the environment. These actions then update the environment’s state in real-time. All these components can be adapted easily to experiment with different social media platforms. For instance, by adjusting specific modules, switching from one platform, such as X, to another like Reddit is possible.
@@ 3.4.2 Herd Effect in Reddit
3.5 Simulating Large-Scale Misinformation Spreading on Platform X using OASIS
In the following subsection, we focus on Platform X (formerly Twitter), analyzing how both true and false information spread among millions of agents. Unlike Reddit, the action space on Twitter includes creating posts, repost, like post, dislike post, follow, create comment, like comment, and dislike comment. Additionally, we shift from the hot-score based recommendation system to an interest-based one.
Experiment Setting
We select four news stories from official sources, covering health, technology, entertainment, and education. We then fabricate four pieces of misinformation that closely resemble the official news. Details are provided in Appendix G.0.1. The experiment includes 196 core agents (with a large number of followers) and 1 million regular agents, as described in Section LABEL:sec:large-scale_agents_generation. e ask the same core user to post both the true and fake versions of each of the four pieces of news. The simulation runs for 60 time steps, with an activation probability of 0.1 for core users and 0.01 for regular users. The experiment is conducted on 24 A100 GPUs within a week.
Misinformation is more influential than the official news.
To investigate the impact of official news and misinformation on the new posts generated by agents, we employ the TF-IDF-based approach (Term Frequency-Inverse Document Frequency Christian et al. (2016)). Specifically, we calculate the cosine similarity between 8 news (including 4 pairs of official and misinformation news) and a simulated set of 733824 posts generated by agents. A similarity threshold of 0.2 was set, with posts exceeding this threshold considered relevant to the target news. We then track the number of posts related to official news and misinformation at each time step, and the results are shown in the Figure 9. As illustrated in the figure, for each topic, the number of posts related to misinformation consistently exceeds those related to official news. In the early stages, the number of posts related to both official news and misinformation increases rapidly due to the smaller number of relevant posts. However, over time, the number of posts declines quickly as posts on other topics gain more traction. Even in the later stages, posts related to misinformation maintain a higher level of activity, suggesting that misinformation has a more sustained influence.
Furthermore, We visualize the new attention relationships formed during the simulation process. We find that these new connections exhibit a certain clustering phenomenon, where users tend to form concentrated clusters, with some forming densely connected central areas and others being more isolated. These subgraphs indicate the existence of complex relational networks, suggesting distinct communities within the overall structure.
Figure 9: TThe figure shows the number of posts related to official news and misinformation at various time steps across different topics. For each topic, posts related to misinformation are more numerous than those related to official news.
Figure 10: Graph of newly established user relationships during the simulation process. Each arrow represents a new follow relationship, and each node represents an agent. Clustering is observable among these new connections.